# Bridging Quantum and Classical Computing in Drug Design: Architecture Principles for Improved Molecule Generation

### v1

@@ Abstract
Hybrid quantum-classical machine learning offers a path to leverage noisy intermediate-scale quantum (NISQ) devices for drug discovery, but optimal model architectures remain unclear. We systematically optimize the quantum-classical bridge architecture for generative adversarial networks (GANs) in molecular discovery using multi-objective Bayesian optimization. Our optimized model (BO-QGAN) significantly improves performance, achieving a 2.27-fold higher Drug Candidate Score (DCS) than prior quantum-hybrid benchmarks and 2.21-fold higher than the classical baseline, using over 60% fewer parameters. Key findings favor layering multiple (3-4) shallow (4-8 qubit) quantum circuits sequentially, while classical architecture shows less sensitivity above a minimum capacity. This work provides the first empirically grounded architectural guidelines for hybrid models, enabling more effective integration of current quantum computers into pharmaceutical research pipelines.
Machine Learning, ICML
@@ 1 Background and Significance
Developing new pharmaceuticals is notoriously lengthy and expensive, often exceeding a decade and billions of dollars (Hughes et al., 2011; Singh et al., 2023). Computational methods offer a path to accelerate this process, but identifying promising candidates within the vast chemical space (estimated at $10^{60}$ molecules) remains a significant hurdle (Bassani & Moro, 2023).
@@ 1.1 Hybrid Quantum-Classical Models: A Pragmatic Approach
Hybrid quantum-classical models offer a pragmatic solution, combining smaller quantum circuits (capturing key quantum features) with larger classical networks (handling scale and decoding). Parameterized quantum circuits, analogous to classical layers, contain tunable parameters and adjustable width (qubits) and depth (layers) (Benedetti et al., 2019). A quantum computation result is “bridged” into the classical network, leveraging quantum expressivity within current hardware limits. We use Generative Adversarial Networks (GANs) as our framework due to established classical (MolGAN) and hybrid (Li et al., 2021) benchmarks, allowing focused analysis of the hybrid architecture itself.
The foundational hybrid model for this task, QGAN-HG (Li et al., 2021), adapted MolGAN (De Cao & Kipf, 2022) and matched its performance with significantly fewer parameters. Subsequent work explored variations like Wasserstein loss and gradient penalty integration (Jain & Ganguly, 2022), different quantum component placements (Kao et al., 2023), and advanced GAN structures like CycleGAN (Anoshin et al., 2024). The potential real-world applicability was highlighted by work generating and experimentally validating novel KRAS inhibitor candidates discovered using quantum-enhanced methods (Ghazi Vakili et al., 2025). While valuable, these studies often focus on framework modifications, overlooking architectural details. To isolate the impact of the quantum-classical interface architecture, we benchmark directly against the original MolGAN and QGAN-HG, anticipating our derived principles may benefit more advanced frameworks later.
Designing the optimal quantum-classical architecture involves complex trade-offs between quantum expressivity, NISQ hardware limitations, and overall model parameter count. Establishing systematic principles to balance these factors for effective information transfer and performance is crucial but currently lacks rigorous investigation.
@@
By establishing the first empirically grounded architectural guidelines, this work demonstrates substantial performance gains (2.3x $DCS$ increase over prior hybrid work) with reduced computational cost ($>$60% fewer parameters than classical baseline), offering a practical path to more effectively integrate NISQ computers into drug discovery.
@@ 2.2.1 Generator Network ($G_{\theta}$)
The generator maps noise $z\in\mathbb{R}^{B\times M}$ (batch size $B$, quantum width $M$) to a molecular graph $\mathcal{G}$. The input $z$ is embedded into a quantum state using angle encoding. This state is transformed by parameterized quantum circuit layers consisting of single-qubit rotations (RY gates) and CNOT-based entangling layers (ring structure). Measurement of Pauli-Z expectation values yields a quantum latent representation ($\mathbb{R}^{B\times M}$). This representation bridges to the classical component: fully connected layers ($\tanh$ activation) output dense tensors $A$ and $X$. Gumbel-Softmax sampling produces discrete graph tensors $\widetilde{A}$ and $\widetilde{X}$. Key optimized hyperparameters are quantum width ($M$), quantum depth (number of circuit layers), classical width (neurons/layer), and classical depth (number of layers).
@@ 2.2.2 Discriminator (Critic) Network
A classical graph neural network distinguishes real vs. generated molecules ($\widetilde{A},\widetilde{X}$), outputting a scalar Wasserstein critic score $r$. It uses graph convolutions and aggregation layers.
@@ 2.2.3 Reward Network
Architecturally similar to the discriminator but with independent parameters, this network predicts the expected reward (desirable chemical properties, Section 2.3) for a generated molecule $\mathcal{G}$, guiding the generator via differentiable gradients (De Cao & Kipf, 2022).
@@ 2.3 Evaluation Metrics
    where 10 is a scaling factor for readability. Individual metrics ($QED,logP,SA$) are normalized to $[0,1]$ (higher is better; note $SA$ normalization reverses standard interpretation).
@@ 2.4 Bayesian Optimization of Bridge Architecture
The search space covered quantum width ($[4,16]$ qubits), quantum depth ($[1,4]$ layers), classical width ($[16,512]$ neurons), and classical depth ($[1,4]$ layers).
@@ 2.6 Experimental Design
Phase 1 involved 100 MOTPE trials optimizing the architecture. Each trial trained a hybrid GAN, using a simulated quantum backend, for up to 5000 epochs with early stopping (no validation $DCS$ improvement for 250 epochs after reaching $FD<12.5$). Phase 2 benchmarked the best $DCS$ architecture discovered during Phase 1 against re-implemented MolGAN (De Cao & Kipf, 2022) and QGAN-HG MR (Li et al., 2021) using identical protocols. Performance metrics were averaged over 30 independent runs (1000 molecules/run). Statistical significance was assessed using t-tests, and effect size was assessed using Cohen’s d. Experiments used Python 3.12, PyTorch 2.4.1, PyG, PennyLane, Optuna, RDKit.
@@ 3.1 Bayesian Optimization Performance
The MOTPE algorithm efficiently navigated the hyperparameter space over 100 trials, progressively identifying architectures with improved $DCS$. The algorithm effectively explored the trade-off between $DCS$ and training time, identifying a Pareto front of optimal configurations. The final model, Bayesian-Optimized QGAN (BO-QGAN), was selected from this front for maximizing $DCS$.
@@ 3.3 Characteristics of Optimized Architecture
The optimization revealed complex, non-linear relationships between model capacity and performance. The best architecture, Bayesian Optimized Quantum GAN (BO-QGAN), uses 3 quantum circuit layers (7 qubits each) feeding into 2 classical layers (227 neurons each) (Table 2). Its superior $DCS$ stems primarily from improved lipophilicity ($logP$) and synthetic accessibility ($SA$) compared to baselines (Table 1). A sample of molecules generated by a model, illustrating typical structures and their corresponding $QED$ scores, is shown in Figure 3.
@@ 3.4 Influence of Quantum Architecture
The quantum component design was critical. Performance peaked within a middle range of total quantum parameters, suggesting a trade-off between expressivity and computational cost/overfitting. Analysis of width vs. depth (Figure 4) revealed that the best models consistently used multiple (3-4) shallow (4-8 qubits) quantum circuits layered sequentially, corresponding to width-to-depth ratios between $\approx 1.0$ and $2.7$. This indicates layering smaller circuits is more effective than single deep/wide circuits for this task.
@@
(c) Comparing the number of qubits used (width) to the $DCS$. The red line and r score show the linear correlation between variables. The green region contains all the configurations with the top 10% of $DCS$.
@@ 3.5 Influence of Classical Architecture
In contrast, the classical architecture showed less sensitivity beyond a minimum capacity threshold (approx. $>$90 neurons per layer). Increasing classical parameters or varying the width-to-depth ratio showed little correlation with improved $DCS$ once this threshold was met (Figure 5). This suggests the classical network primarily needs sufficient capacity to process the quantum output, with diminishing returns from further size increases.
@@ 3.6 Interplay between Quantum and Classical Architecture
Analysis revealed no strong interplay between optimal quantum and classical sizes. Figure 6 visualizes the achieved $DCS$ across the search space defined by quantum and classical parameter counts. The landscape appears noisy, without clear ridges which would indicate that certain quantum sizes work best only with specific classical sizes. The correlation between the ratio of quantum-to-classical parameters and $DCS$ was weak ($r=-0.20$). Optimal performance seems to depend on finding suitable configurations within each component independently, rather than a specific balance between them.
@@ 3.7 Computational Efficiency
The optimization process successfully balanced performance and computational cost. Figure 7 plots the final $DCS$ achieved against the total training time for each of the 100 trials. No significant correlation was found between longer training times and better final $DCS$ ($r=-0.01$). The best model, BO-QGAN, trained efficiently ($<$17 minutes on test hardware, 2140 iterations). Training time in simulation was primarily driven by quantum circuit width (qubits), consistent with the exponential scaling of simulation cost with $N$. The finding that shallower, narrower circuits perform best aligns well with computational tractability.
@@ 3.8 Validation on Real Quantum Hardware
We generated a batch of 10 molecular structures using this high-validity configuration executed on IBM’s 127-qubit QPU ’Sherbrooke’ via the cloud platform. For direct comparison, a batch of 10 molecules was generated using the same model configuration executed on a classical quantum simulator (PennyLane’s ‘default.qubit‘). The resulting molecules from both backends were evaluated using the standard druglikeness metrics ($QED$, $logP$, $SA$, $DCS$).
@@ 4.2 Design Principles for Quantum Circuits
Based on the optimization results, particularly the analysis of the influence of quantum architecture (Section 3.4), we can propose practical design guidelines for hybrid generative models in this domain. For the QM9 dataset, characterized by small molecules, the best-performing models consistently favored multiple (3-4), sequentially layered, shallow (4-8 qubits) quantum circuits over fewer, wider or deeper ones. This suggests a core principle: when designing the quantum component, prioritize layering relatively shallow circuits sequentially.
The potential reasons for the effectiveness of this layered, shallow approach are multifaceted. It might relate to improved trainability on NISQ devices; shallower circuits are generally less susceptible to noise accumulation and may mitigate the risk of encountering barren plateaus in the optimization landscape (McClean et al., 2018). Furthermore, layering circuits could allow the model to learn hierarchical representations, with earlier layers capturing basic features and later layers combining them into more complex patterns, analogous to deep classical networks but within the quantum domain. This approach also aligns well with computational efficiency, especially for simulation, where cost scales exponentially with qubit count (width) but only polynomially with gate depth (layers). Therefore, when allocating limited quantum resources, prioritizing adding circuit layers over adding qubits appears to be a more cost-effective strategy for improving performance on this type of task.
Regarding the classical component, the principle appears simpler: ensure sufficient capacity (roughly $>$90 neurons per layer in our experiments) to process the information from the quantum bridge, after which further increases yield diminishing returns (Section 3.5). These empirically derived principles offer valuable, quantitative starting points for designing effective and resource-aware quantum-classical hybrid models in the NISQ era.
@@ 4.3 Limitations and Mitigation
  • QM9 Dataset Scope: QM9 is a standard benchmark in this field, but it contains small, simple molecules. Generalization to larger, more diverse drug candidates (e.g., from ZINC, ChEMBL) requires further investigation.
  • MolGAN Baseline Framework: Performance gains are relative to the MolGAN framework. A simple model was chosen to help isolate the effect of the architecture optimization. Applying the optimization methodology and derived principles to state-of-the-art generative models (e.g., diffusion) is a logical next step.
  • $DCS$ Metric and Optimization Scope: The $DCS$ metric is one way to quantify quality; results might differ with other objectives. Optimization focused on width/depth; exploring gates, embedding, or measurement could yield further gains.
@@ 4.4 Future Research Directions
Future work should prioritize extensive validation on diverse NISQ hardware platforms with larger sample sizes and error mitigation. Applying the optimization methodology to larger datasets (ZINC, ChEMBL) and more advanced generative frameworks (diffusion, flow-matching) is crucial for assessing scalability and generalization. Further investigation into optimizing finer architectural details (gates, embedding, measurement) and adapting the framework for target-specific objectives (e.g., binding affinity) are also promising directions.

### v2

@@ Abstract
Hybrid quantum-classical machine learning offers a path to leverage noisy intermediate-scale quantum (NISQ) devices for drug discovery, but optimal model architectures remain unclear. We systematically optimize the quantum-classical bridge architecture of generative adversarial networks (GANs) for molecule discovery using multi-objective Bayesian optimization. Our optimized model (BO-QGAN) significantly improves performance, achieving a 2.27-fold higher Drug Candidate Score (DCS) than prior quantum-hybrid benchmarks and 2.21-fold higher than the classical baseline, while reducing parameter count by more than 60%. Key findings favor layering multiple (3-4) shallow (4-8 qubit) quantum circuits sequentially, while classical architecture shows less sensitivity above a minimum capacity. This work provides the first empirically-grounded architectural guidelines for hybrid models, enabling more effective integration of current quantum computers into pharmaceutical research pipelines.
Machine Learning, ICML, Drug Discovery
@@ 1 Background and Significance
Developing new pharmaceuticals is a notoriously long and expensive process, often exceeding a decade and costing billions of dollars (Hughes et al., 2011; Singh et al., 2023). Computational methods offer a path to accelerate this process, but identifying promising candidates within the vast chemical space (estimated at $10^{60}$ molecules) remains a significant hurdle (Bassani & Moro, 2023).
@@ 1.1 Hybrid Quantum-Classical Models: A Pragmatic Approach
Hybrid quantum-classical models offer a pragmatic solution, combining smaller quantum circuits (capturing key quantum features) with larger classical networks (handling scale and decoding). Parameterized quantum circuits, which are analogous to classical layers, contain tunable parameters and have adjustable width (qubits) and depth (layers) (Benedetti et al., 2019). A quantum computation result is bridged into the classical network, leveraging quantum expressivity within current hardware limits. We use Generative Adversarial Networks (GANs) as our framework due to established classical (MolGAN) and hybrid (Li et al., 2021) benchmarks, allowing focused analysis of the hybrid architecture itself.
The original hybrid model for this task, QGAN-HG (Li et al., 2021), adapted MolGAN (De Cao & Kipf, 2022) and matched its performance with significantly fewer parameters. Subsequent work explored variations like Wasserstein loss and gradient penalty integration (Jain & Ganguly, 2022), different quantum component placements (Kao et al., 2023), and advanced GAN structures like CycleGAN (Anoshin et al., 2024). The potential real-world applicability was highlighted by work generating and experimentally validating novel KRAS inhibitor candidates discovered using quantum-enhanced methods (Ghazi Vakili et al., 2025). While valuable, these studies often focus on framework modifications, overlooking architectural details. To isolate the impact of the quantum-classical interface architecture, we benchmark directly against the original MolGAN and QGAN-HG, anticipating that our derived principles may benefit more advanced frameworks later.
Designing the optimal quantum-classical architecture involves complex trade-offs between quantum expressivity, NISQ hardware limitations, and overall model parameter count. Establishing systematic principles to balance these factors for effective information transfer and performance is crucial, but currently lacks rigorous investigation.
@@
By establishing the first empirically grounded architectural guidelines, this work demonstrates substantial performance gains (2.3x $DCS$ increase over previous hybrid work) with reduced computational cost ($>$60% fewer parameters than the classical baseline), offering a practical path to integrate NISQ computers more effectively into drug discovery.
@@ 2.2.1 Generator Network ($G_{\theta}$)
The generator maps noise $z\in\mathbb{R}^{B\times M}$ (batch size $B$, quantum width $M$) to a molecular graph $\mathcal{G}$. The input $z$ is embedded into a quantum state using angle encoding. This state is transformed by parameterized quantum circuit layers consisting of single-qubit rotations (RY gates) and CNOT-based entangling layers (with a ring structure). Measurement of Pauli-Z expectation values yields a quantum latent representation ($\mathbb{R}^{B\times M}$). This representation is transferred to the classical component, where fully connected layers (with $\tanh$ activations) output dense tensors $A$ and $X$. Gumbel-Softmax sampling produces discrete graph tensors $\widetilde{A}$ and $\widetilde{X}$. Key optimized hyperparameters are quantum width ($M$), quantum depth (number of circuit layers), classical width (neurons/layer), and classical depth (number of layers).
@@ 2.2.2 Discriminator (Critic) Network
The discriminator is a classical graph neural network, which distinguishes real vs. generated molecules ($\widetilde{A},\widetilde{X}$), outputting a scalar Wasserstein critic score $r$. It uses graph convolutions and aggregation layers.
@@ 2.2.3 Reward Network
Architecturally similar to the discriminator but with independent parameters, the reward network predicts the expected reward (desirable chemical properties, Section 2.3) for a generated molecule $\mathcal{G}$, guiding the generator via differentiable gradients (De Cao & Kipf, 2022).
@@ 2.3 Evaluation Metrics
    where 10 is a scaling factor for readability. Individual metrics ($QED,logP,SA$) are normalized to [0,1][0,1][ 0 , 1 ] (higher is better; note $SA$ normalization reverses standard interpretation).
@@ 2.4 Bayesian Optimization of Bridge Architecture
The search space covered quantum width ([4,16][4,16][ 4 , 16 ] qubits), quantum depth ([1,4][1,4][ 1 , 4 ] layers), classical width ([16,512][16,512][ 16 , 512 ] neurons), and classical depth ([1,4][1,4][ 1 , 4 ] layers).
@@ 2.6 Experimental Design
Phase 1 involved 100 MOTPE trials that optimized the architecture. Each trial trained a hybrid GAN, using a simulated quantum backend for up to 5000 epochs with early stopping (no validation $DCS$ improvement for 250 epochs after reaching $FD<12.5$). Phase 2 benchmarked the best $DCS$ architecture discovered during Phase 1 against re-implemented MolGAN (De Cao & Kipf, 2022) and QGAN-HG MR (Li et al., 2021) using identical protocols. Performance metrics were averaged over 30 independent runs (1000 molecules/run). Statistical significance was assessed using t-tests, and effect size was assessed using Cohen’s d. Experiments used Python 3.12, PyTorch 2.4.1, PyTorch Geometric, PennyLane, Optuna, RDKit (Paszke et al., 2019; Fey & Lenssen, 2019; Bergholm et al., 2022; Akiba et al., 2019; Landrum et al., 2024).
The results of a hybrid quantum-classical model discovered by the optimization process were then validated on real quantum hardware.
@@ 3.1 Bayesian Optimization Performance
The MOTPE algorithm efficiently navigated the hyperparameter space in the 100 trials, progressively identifying improved architectures $DCS$. The algorithm effectively explored the trade-off between $DCS$ and training time, identifying a Pareto front of optimal configurations. The final model, which we call Bayesian-Optimized Quantum GAN (BO-QGAN), was selected from this front to maximize $DCS$.
@@ 3.3 Characteristics of Optimized Architecture
The optimization revealed complex, non-linear relationships between model capacity and performance. The best architecture, Bayesian-Optimized Quantum GAN (BO-QGAN), uses 3 quantum circuit layers (7 qubits each) feeding into 2 classical layers (227 neurons each) (Table 2). Its superior $DCS$ is mainly due to improved lipophilicity ($logP$) and synthetic accessibility ($SA$) compared to baselines (Table 1). A sample of molecules generated by a model and their corresponding $QED$ scores is shown in Figure 3.
@@ 3.4 Influence of Quantum Architecture
The quantum component design was critical. Performance peaked within a middle range of total quantum parameters, suggesting a trade-off between expressivity and computational cost/overfitting. Analysis of width vs. depth (Figure 4) revealed that the best models consistently used multiple (3-4) shallow (4-8 qubits) quantum circuits layered sequentially, corresponding to width-to-depth ratios between 1.0$\approx$ 1.0≈ 1.0 and 2.72.72.7. This indicates that layering more smaller circuits is more effective than using fewer, wider circuits for this task.
@@
(c) Comparing the number of qubits used (width) to the $DCS$. The red line and r score show the linear correlation between variables. The green region contains all the configurations within the top 10% of $DCS$.
@@ 3.5 Influence of Classical Architecture
In contrast, the classical architecture showed less sensitivity beyond a minimum capacity threshold (approx. $>$90 neurons per layer). Increasing classical parameters or varying the width-to-depth ratio showed little correlation with improved $DCS$ once this threshold was met (Figure 5). This suggests that the classical network primarily needs sufficient capacity to process the quantum output, with diminishing returns from further size increases.
@@ 3.6 Interplay between Quantum and Classical Architecture
Analysis revealed no strong interplay between optimal quantum and classical network sizes. Figure 6 visualizes the $DCS$ achieved across the search space (defined by quantum and classical parameter counts). The landscape appears noisy, without clear ridges which would indicate that certain quantum sizes work best only with specific classical sizes. The correlation between the ratio of quantum-to-classical parameters and $DCS$ was weak ($r=-0.20$). Optimal performance seems to depend on finding suitable configurations within each component independently, rather than a specific balance between them.
@@ 3.7 Computational Efficiency
The optimization process successfully balanced performance and computational cost. Figure 7 plots the final $DCS$ achieved against the total training time for each of the 100 trials. No significant correlation was found between longer training times and better final $DCS$ ($r=-0.01$). The best model, BO-QGAN, trained efficiently ($<$17 minutes on test hardware, using 2140 training iterations). Training time with the simulated backend was primarily driven by quantum circuit width (qubits), consistent with the exponential scaling of simulation cost with $N$. The finding that shallower, narrower circuits perform best aligns well with computational tractability.
@@ 3.8 Validation on Real Quantum Hardware
We generated a batch of 10 molecular structures using this high-validity configuration executed on IBM’s 127-qubit QPU ‘Sherbrooke’ via the cloud platform. For direct comparison, a batch of 10 molecules was generated using the same model configuration executed on a classical quantum simulator (PennyLane’s ‘default.qubit‘). The resulting molecules from both backends were evaluated using the standard druglikeness metrics ($QED$, $logP$, $SA$, $DCS$).
@@ 4.2 Design Principles for Quantum Circuits
Based on the optimization results, particularly the analysis of the influence of quantum architecture (Section 3.4), we can propose practical design guidelines for hybrid generative models in this domain. For the QM9 dataset, characterized by small molecules, the best-performing models consistently favored multiple (3-4), sequentially layered, shallow (4-8 qubits) quantum circuits over fewer wider or deeper ones. This suggests a core principle: when designing the quantum component, prioritize layering relatively shallow circuits sequentially.
The potential reasons for the effectiveness of this layered, shallow approach are multifaceted. It may relate to improved trainability on NISQ devices; shallower circuits are generally less susceptible to noise accumulation and may mitigate the risk of encountering barren plateaus in the optimization landscape (McClean et al., 2018). Furthermore, layering circuits could allow the model to learn hierarchical representations, with earlier layers capturing basic features and later layers combining them into more complex patterns, analogous to deep classical networks but within the quantum domain. This approach also aligns well with computational efficiency, especially for simulation, where cost scales exponentially with qubit count (width) but only polynomially with gate depth (layers). Therefore, when allocating limited quantum resources, prioritizing adding circuit layers over adding qubits appears to be a more cost-effective strategy for improving performance on this type of task.
Regarding the classical component, the principle appears simpler: ensure sufficient capacity (roughly $>$90 neurons per layer in our experiments) to process the information from the quantum circuits, after which further increases yield diminishing returns (Section 3.5). These empirically derived principles offer valuable, quantitative starting points for designing effective and resource-aware quantum-classical hybrid models in the NISQ era.
@@ 4.3 Limitations and Mitigation
  • QM9 Dataset Scope: QM9 is a standard benchmark in this field, but it contains small, simple molecules. Generalizability to larger, more diverse drug candidates (e.g., from ZINC or ChEMBL) requires further investigation.
  • MolGAN Baseline Framework: Performance gains are relative to the MolGAN framework. A simple model was chosen to help isolate the effect of the architecture optimization. Applying the optimization methodology and derived principles to state-of-the-art generative models (e.g. diffusion) is a logical next step.
  • $DCS$ Metric and Optimization Scope: The $DCS$ metric is one way to quantify quality; results might differ with other objectives. The optimization experiment focused on network width and depth; exploring different gates, embeddings, or measurements could yield further gains.
@@ 4.4 Future Research Directions
Future work should prioritize extensive validation on diverse NISQ hardware platforms with larger sample sizes and error mitigation. Applying the optimization methodology to larger datasets (ZINC, ChEMBL) and more advanced generative frameworks (diffusion, flow-matching) is crucial for assessing scalability and generalizability. Further investigation into optimizing finer architectural details (gates, embedding, measurement) and adapting the framework for target-specific objectives (binding affinity) are also promising directions.